# Linear Attention Sequence Parallelism

### v1

@@ Abstract
Sequence Parallel (SP) serves as a prevalent strategy to handle long sequences that exceed the memory limit of a single GPU. However, existing SP methods do not take advantage of linear attention features, resulting in sub-optimal parallelism efficiency and usability for linear attention-based language models. In this paper, we introduce Linear Attention Sequence Parallel (LASP), an efficient SP method tailored to linear attention-based language models. Specifically, we design an efficient point-to-point communication mechanism to leverage the right-product kernel trick of linear attention, which sharply decreases the communication overhead of SP. We also enhance the practical efficiency of LASP by performing kernel fusion and intermediate state caching, making the implementation of LASP hardware-friendly on GPU clusters. Furthermore, we meticulously ensure the compatibility of sequence-level LASP with all types of batch-level data parallel methods, which is vital for distributed training on large clusters with long sequences and large batches. We conduct extensive experiments on two linear attention-based models with varying sequence lengths and GPU cluster sizes. LASP scales sequence length up to 4096K using 128 A100 80G GPUs on 1B models, which is 8$\times$ longer than existing SP methods while being significantly faster. The code is available at https://github.com/OpenNLPLab/LASP.
Linear attention, sequence parallelism, unlimited sequence length, large language model
@@ 1 Introduction
Recently, linear attention-based models (Qin et al., 2024a, 2022a; Choromanski et al., 2020) are becoming increasingly popular due to their faster processing speed and comparable language modeling performance to vanilla Softmax transformers (Vaswani et al., 2017). Meanwhile, the growing size of large language models (LLMs) (Touvron et al., 2023a, b; Tang et al., 2024) and longer sequence lengths exert significant strain on contemporary GPU hardware because the memory of a single GPU confines the maximum sequence length of a language model. To tackle this issue, the Sequence Parallelism (SP) techniques (Li et al., 2022; Korthikanti et al., 2022) are often utilized to divide a long sequence into several sub-sequences and train them on multiple GPUs separately. Unfortunately, existing SP methods do not fully leverage the benefits of linear attention features, leading to suboptimal parallelism efficiency and usability.
In this paper, we present the Linear Attention Sequence Parallel (LASP) technique for efficient sequence parallelism on linear transformers. Our approach includes a sophisticated communication mechanism based on point-to-point (P2P) communication for exchanging intermediate states during forward and backward passes among GPUs within a node or across multiple nodes. This design maximizes the utilization of right-product kernel tricks (Katharopoulos et al., 2020; Qin et al., 2023b, 2024b) in linear attention. Notably, our technique is independent of attention heads partitioning, which allows it to be applied to models with varying numbers or styles of attention heads, such as multi-head, multi-query, and grouped-query attentions. This flexibility exceeds the capabilities of existing SP methods in Megatron-LM (Shoeybi et al., 2019; Korthikanti et al., 2022) or DeepSpeed (Jacobs et al., 2023).
Our implementation of LASP incorporates system engineering optimizations such as kernel fusion and KV State caching, resulting in significantly enhanced execution efficiency. Furthermore, we have taken care in ensuring compatibility of LASP with various (sharded) distributed data-parallel (DDP) (Li et al., 2020; Sun et al., 2024) training methods during the implementation, which we refer to as the data-sequence hybrid parallelism. Through extensive experiments with models of different parameters, cluster sizes, and sequence lengths, we demonstrate LASP’s impressive performance and efficiency when used with these DDP instances. Specifically, LASP is significantly faster than existing SP methods and can extend sequence length 8$\times$ longer under the same hardware constraints.
@@
  • A new SP strategy that is tailored to linear attention. Enabling linear attention-based models to scale for long sequences without being limited by a single GPU.
  • Sequence length-independent communication overhead. Our elegant communication mechanism leverages right-product kernel trick of linear attention to ensure that the exchanging of linear attention intermediate states is sequence length-independent.
  • GPU friendly implementation. We optimize LASP’s execution on GPUs through meticulous system engineering, including kernel fusion and KV State caching.
  • Data-parallel compatibility. LASP is compatible with all batch-level DDP methods, such as PyTorch/Legacy DDP, FSDP, and ZeRO-series optimizers.
@@ 2.1 Preliminary
$\small\mathbf{O}=\mathrm{Softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d}\odot\mathbf{M})\mathbf{V},$ (1)
@@
Linear Attention. Linear attention is originally proposed in (Katharopoulos et al., 2020), with the elimination of Softmax operation (Vaswani et al., 2017). Qin et al. (2022a, 2023a) propose to replace the Softmax operation with a normalization operation $\mathrm{Norm(\cdot)}$, which turns to a concise formulation as:
$\small\mathbf{O}=\mathrm{Norm}((\mathbf{Q}\mathbf{K}^{\top}\odot\mathbf{M})\mathbf{V}).$ (2)
When considering bidirectional tasks, the above formulation can be simplified as $\mathbf{O}=\mathrm{Norm}((\mathbf{Q}\mathbf{K}^{\top})\mathbf{V})$. Then by performing the associativity property of matrix products, it can be mathematically equivalently transformed into a right-product version:
$\small\mathbf{O}=\mathrm{Norm}(\mathbf{Q}(\mathbf{K}^{\top}\mathbf{V})).$ (3)
@@
Figure 1: Visualization of LASP. We illustrate a typical linear transformer layer to explain the mechanism of LASP. $\mathbf{X}$ as an input sequence is partitioned into multiple sub-sequence chunks, which are then fed into different model copies on discrete devices. $g$ represents conjugate communication operations in forward and backward passes. In forward pass, $g$ is Send &\&& Recv from device $i$ to $(i+1)$. While in backward pass, $g$ is Send &\&& Recv from device $(i+1)$ to $i$. The communication operations exchange intermediate states $\mathbf{KV}$ and $\mathbf{dKV}$ during forward and backward passes to insure the performance of SP.
@@ 2.2 LASP
LASP tiles sequence over the cluster. Follow the thought-of-tiling, LASP partitions the input sequences into multiple sub-sequence chunks, distributing these chunks individually across different GPUs. For linear attention in a casual setting, in order to fully exploit the advantage of right-product in linear attention, we categorize the attention computation for sub-sequences into two distinct types: intra-chunks and inter-chunks. Intra-chunks involve conventional attention computation, while inter-chunks leverage the kernel tricks associated with linear attention’s right-product. Further details regarding the intricate mechanisms of LASP in data distribution, forward pass, and backward pass are expounded upon below. A visualization of LASP is presented in Fig. 1.
Data Distribution.
For an input sequence of length $N$, we establish its embedding space representation denoted as $\mathbf{X}\in\mathbb{R}^{N\times d}$ with a feature dimension of $d$. In the LASP framework, $\mathbf{X}$ is evenly partitioned into $T$ chunks, where $T$ is called the sequence parallel size, which must be divisible by the distributed world size $W$. These segmented data chunks are subsequently assigned to the respective GPUs. It is essential to note that different sequence parallel groups receive dissimilar data batches. However, within the same group, all data chunks originate from an identical batch of data. A comprehensive depiction of the data distribution process in LASP is provided in Algorithm 1. Additionally, an illustrative example of data distribution in LASP is presented in Fig. 2, considering a node with 8 GPUs and the partitioning of 2 sequences into 4 sub-sequence chunks.
@@
Figure 2: Example of LASP Data Distribution. In this illustration, the distributed world size is characterized by $W=8$, the sequence parallel size by $T=4$, the number of sequence parallel groups by $G=2$, and the sequence parallel source rank list by $R_{{src}}=[0,4]$. For the first batch Seq0, the input sequence ${\mathbf{X}}$ undergoes partitioning into $T$ chunks $\{\mathbf{X}_{1},\mathbf{X}_{2},...\mathbf{X}_{T}\}$ along the sequence dimension, subsequently transmitted to the first rank in the first SP group denoted as SP-Group0, which corresponds to global rank 0. The data chunks on global rank 0 are then scattered to global ranks 0, 1, 2, 3 within SP-Group0, where each rank only retains a single chunk. The subsequent batch Seq1 follows a similar allocation procedure, being assigned to global ranks 4, 5, 6, 7 within SP-Group1.
@@
13:    Save $\mathbf{KV}_{{t-1}}$ as $\mathbf{KV}_{i}$ on rank $i$ for backward computation.
@@
15:    Compute $\mathbf{O}_{t}=\mathbf{O}_{\mathrm{t,intra}}+\mathbf{O}_{\mathrm{t,inter}}$ as $\mathbf{O}$ of $t$-th chunk.
@@
Forward Pass.
To streamline derivations, the $\mathrm{Norm}(\cdot)$ operator in Eq. (2) is temporarily omitted. Additionally, we consider a normal case where $W=T$, indicating $G=W/T=1$. In this scenario, GPU with rank 00 consolidates all split sub-sequences in a batch, subsequently distributing them to all GPUs across the entire distributed world. It is noteworthy that the scenario where the sequence parallel size is not equal to world size is discussed in 2.5.
We first define two symbols $\mathbf{kv}$ and $\mathbf{KV}$ which are important in the right-product computation procedure of linear attention. Without loss of generality, we add $\lambda$ as the decay rate in linear attention with casual mask, choosing $\lambda=1$ yields the ordinary linear attention. In the forward pass of linear attention computation with casual mask, the $s$-th output can be formulated as
$\small\mathbf{o}_{s}^{\top}=\mathbf{q}_{s}^{\top}\sum_{i\leq s}\lambda^{s-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}.$ (4)
Rewrite in a recursive form (Katharopoulos et al., 2020), we have
$\mathbf{kv}_{0}$ $=0\in\mathbb{R}^{d\times d},$ (5)
$\mathbf{kv}_{s}$ $=\lambda\mathbf{kv}_{s-1}+\mathbf{k}_{s}\mathbf{v}_{s}^{\top},$
$\mathbf{o}_{s}^{\top}$ $=\mathbf{q}_{s}^{\top}(\mathbf{kv}_{s}),$
@@
$\small\textstyle\mathbf{kv}_{s}=\sum_{i\leq s}\lambda^{s-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}$ (6)
is the activation state in the forward computation of linear attention with $s$-th input.
In the sequence parallelism scenario, given data chunk $\mathbf{X}_{t}$ on rank $i$, the query, key and value corresponding to $\mathbf{X}_{t}$ is $\mathbf{Q}_{t}=\mathbf{X}_{t}\mathbf{W}_{Q}$, $\mathbf{K}_{t}=\mathbf{X}_{t}\mathbf{W}_{K}$, $\mathbf{V}_{t}=\mathbf{X}_{t}\mathbf{W}_{V}$. Note that we assume $T=W$ here, their indices are thus equivalent, i.e., $t=i$. The output within the $t$-th chunk can be calculated as
$\small\mathbf{O}_{\mathrm{t,intra}}=[(\mathbf{Q}_{t}\mathbf{K}_{t}^{\top})\odot\mathbf{M}]\mathbf{V}_{t}.$ (7)
@@
$\mathbf{o}_{s+C}^{\top}=\mathbf{q}_{s+C}^{\top}\sum_{i\leq s+C}\lambda^{s+C-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}$ (8)
$=$ $\mathbf{q}_{s+C}^{\top}\sum_{i=C+1}^{C+s}\lambda^{s+C-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}+\lambda^{s}\mathbf{q}_{s+C}^{\top}\sum_{i\leq C}\lambda^{C-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}.$
The first part (before the plus sign) in Eq. (8) corresponds to the computation on intra-chunk, and the second part (after the plus sign) corresponds to the computation on inter-chunk. In sequence parallelism scenario, Eq. (8) can be rewritten in the chunk form as follows:
$\small\mathbf{O}_{\mathrm{t,inter}}=\bm{\Lambda}\mathbf{Q}_{t}\mathbf{K}\mathbf{V}_{t-1},$ (9)
where
$\small\mathbf{KV}_{t}=\mathbf{kv}_{tC}.$ (10)
It is worth noting that the calculation of the inter part for the $t$-th chunk depends on the activation state of previous $(t-1)$ chunk, i.e., $\mathbf{KV}_{t-1}$, which is calculated on rank $(i-1)$. Thus a P2P communication operation Recv should be performed to pull $\mathbf{KV}_{t-1}$ from rank $(i-1)$ to rank $i$. Then the activation state $\mathbf{KV}_{t}$ should be updated for subsequent inter-chunk attention computation at $(t+1)$-th chunk. The update rule of $\mathbf{KV}_{t}$ at $t$-th chunk is
$\mathbf{KV}_{t}$ $=\sum_{s\leq tC}\lambda^{tC-s}\mathbf{k}_{s}\mathbf{v}_{s}^{\top}$ (11)
$=\lambda^{C}\!\!\!\!\sum_{s\leq(t-1)C}\!\!\!\!\lambda^{(t-1)C-s}\mathbf{k}_{s}\mathbf{v}_{s}^{\top}+\!\!\!\!\!\!\sum_{s=(t-1)C+1}^{tC}\!\!\!\!\!\!\lambda^{tC-s}\mathbf{k}_{s}\mathbf{v}_{s}^{\top}$
$=\lambda^{C}\mathbf{KV}_{t-1}+\left(\mathrm{diag}\{\lambda^{C-1},\ldots,1\}\mathbf{K}_{t}\right)^{\top}\mathbf{V}_{t}$
$=\lambda^{C}\mathbf{KV}_{t-1}+\left(\lambda^{C}\mathbf{\Lambda}^{-1}\mathbf{K}_{t}\right)^{\top}\mathbf{V}_{t}.$

### v3

@@ Abstract
Sequence parallelism (SP) serves as a prevalent strategy to handle long sequences that exceed the memory limit of a single device. However, for linear sequence modeling methods like linear attention, existing SP approaches do not take advantage of their right-product-first feature, resulting in sub-optimal communication efficiency and usability. In this paper, we introduce Linear Attention Sequence Parallelism (LASP), an efficient SP approach designed for linear attention-based transformer models. Specifically, we design an efficient point-to-point ring-style communication mechanism to leverage the right-product kernel trick of linear attention, which sharply decreases the communication overhead, comparing with existing SP methods. We enhance the computation efficiency of LASP by performing kernel fusion and intermediate state caching, making the implementation of LASP hardware-friendly on GPUs. Furthermore, we meticulously ensure the compatibility of sequence-level LASP with all types of batch-level data parallel methods, which is vital for distributed training on large clusters with very-long sequences. We also discuss the generalization of LASP on other linear sequence modeling methods. Extensive experiments on linear attention-based models are conducted with varying sequence lengths from 2K to 4096K. LASP scales sequence length up to 4096K on 128 GPUs, which is 8$\times$ longer than existing SP methods. Code is available at: https://github.com/OpenNLPLab/LASP.
@@ 1 Introduction
Linear sequence modeling methods (Katharopoulos et al., 2020; Choromanski et al., 2022; Sun et al., 2025a) including linear attention (Qin et al., 2024d), state space models (Dao & Gu, 2024) and linear RNN (Qin et al., 2024e), are becoming increasingly popular due to their faster training and inference speed and comparable modeling performance to vanilla Softmax attention-based transformer models (Vaswani et al., 2017; Zeng et al., 2022; Touvron et al., 2023a; b; Team, 2023). The hybrid architecture, which interleaves Softmax attention and linear attention Transformer layers, has proven to be an effective balance between their respective strengths. This approach has been successfully implemented in large-scale commercial models such as Minimax-01 (Li et al., 2025) and Tencent Hunyuan Turbo-S (Tencent, 2025), as well as in smaller-scale hybrid models like Samba (Ren et al., 2024), Jamba (Lieber et al., 2024).
As the size of large language models (LLMs) increases and sequence lengths extend, the capacity limitations of single GPU’s memory become a significant challenge, constraining the maximum sequence length manageable by a large model. To address this, Sequence Parallelism (SP) techniques (Li et al., 2022; Korthikanti et al., 2022) are employed, which partition a long sequence into multiple sub-sequences to be processed on separate devices. However, current implementations of SP methods do not fully exploit the right-product advantages of linear-complexity attention mechanisms Qin et al. (2024b). This results in less than optimal parallelism efficiency and reduced usability on linear sequence modeling methods.
In this paper, we present Linear Attention Sequence Parallelism (LASP) approach for efficient SP on models with linear sequence modeling. Our approach takes linear attention (Katharopoulos et al., 2020) as an instance to design a sophisticated point-to-point (P2P) ring-style communication mechanism during both forward and backward among devices within a node or across multiple nodes. This design maximizes the utilization of right-product kernel tricks in linear attention, by only exchanging one single intermediate state instead of both of key and value states in other counterparts. Notably, our approach is independent of attention heads partitioning, allowing it to be applied to models with varying numbers or styles of attention heads, such as multi-head, multi-query, and grouped-query attentions. This flexibility exceeds the capabilities of existing SP methods in Megatron-LM (Shoeybi et al., 2019; Korthikanti et al., 2022) or DeepSpeed (Jacobs et al., 2023).
Our implementation of LASP incorporates system engineering optimizations such as kernel fusion and KV State caching, resulting in significantly enhanced execution efficiency. Furthermore, we have taken great care in ensuring compatibility of LASP with various (sharded) distributed data-parallel (DDP) (Li et al., 2020) training methods during the implementation, which we refer to as the data-sequence hybrid parallelism. Through extensive experiments with linear transformer models of different parameter numbers, cluster sizes, and sequence lengths, we demonstrate the performance and efficiency of LASP when used with different DDP instances. Specifically, LASP can extend sequence length up to 4096K on 128 GPUs, which is 8$\times$ longer than existing SP methods.
@@
  • A new SP approach called LASP that is designed for linear sequence modeling methods. LASP is able to perform sequence-level distributed training on 8$\times$ longer sequences than existing SP methods while being significantly faster.
  • Sequence length-independent communication overhead. Our proposed P2P ring-style communication strategy leverages right-product kernel trick of linear attention to ensure that the exchanging of linear attention intermediate states is sequence length-independent.
  • GPU friendly implementation. We optimize the execution of LASP on GPU hardware through meticulous system engineering, including kernel fusion and KV State caching.
  • Data-parallel compatibility. LASP is compatible with all batch-level DDP methods, including PyTorch/Legacy DDP, FSDP, and ZeRO-series optimizers.
2 Related Work
Linear Attention. Linear Transformer models bypass the use of Softmax attention by adopting various approximation methods  (Katharopoulos et al., 2020; Peng et al., 2021; Qin et al., 2022a; Shen et al., 2024) instead. The central concept involves using the "kernel trick" to speed up the calculation of the attention matrix, specifically by multiplying keys and values before tackling the computationally intensive $n\times n$ matrix multiplication (Sun et al., 2025b). For instance,  Katharopoulos et al. (2020) use $1+\mathrm{elu}$ activation function,  Qin et al. (2022b) utilizes the cosine function to imitate Softmax characteristics. TransNormerLLM (Qin et al., 2024a) proposes Lightning Attention to accelerate linear attention via optimized IO operations, while Lightning Attention-2 (Qin et al., 2024c) enhances efficiency by separately processing inter- and intra-block computations. RetNet (Sun et al., 2023) integrates retention with attention for parallel training and linear-time inference. GLA (Yang et al., 2023) introduces data-independent gating and a hardware-efficient training algorithm. DeltaNet (Schlag et al., 2021) and its parallelized version (Yang et al., 2024b) improve long-context performance using a delta-rule update. GSA (Zhang et al., 2024), inspired by GLA, incorporates bounded-memory slot control to enhance recall-heavy tasks.
Despite significant research progress, the adoption of linear attention and similar linear sequence modeling techniques in commercial large-scale models remains limited. However, some companies have started exploring their use. For example, Minimax-01 incorporates Lightning Attention (Li et al., 2025), a variant of linear attention; Tencent’s Hunyuan Turbo-S employs Mamba2 (Tencent, 2025), a variant of state-space model ; and Together.AI integrates StripedHyena (Poli et al., 2023), a long convolution model.
Sequence Parallelism. SP as a widely used method to train long sequences has been integrated into many large model training frameworks, including Megatron-LM, DeepSpeed, and Colossal-AI. Megatron-LM (Shoeybi et al., 2019) implements SP along with model (tensor) parallelism (MP) to perform large matrix multiplications on GPUs. However, MP partitions the attention heads, which limits the maximum parallelism degree to be less than the number of attention heads. DeepSpeed-Ulysses (Jacobs et al., 2023) uses an all-to-all communication primitive to reduce communication volume, but also partitions attention heads and faces similar issues as Megatron-LM.
@@ 3.1 Preliminary
$\mathbf{O}=\mathrm{Softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d}\odot\mathbf{M})\mathbf{V},$ (1)
@@
Linear Attention. Linear attention is originally proposed in (Katharopoulos et al., 2020), with the elimination of Softmax operation (Vaswani et al., 2017). Qin et al. (2022a; 2024a) propose to replace the Softmax operation with a normalization operation $\mathrm{Norm(\cdot)}$, which turns to
$\mathbf{O}=\mathrm{Norm}((\mathbf{Q}\mathbf{K}^{\top}\odot\mathbf{M})\mathbf{V}).$ (2)
When considering bidirectional tasks, the above formulation can be simplified as
$\mathbf{O}=\mathrm{Norm}((\mathbf{Q}\mathbf{K}^{\top})\mathbf{V}).$
Then by performing the associativity property of matrix products, it can be mathematically equivalently transformed into a right-product version:
$\mathbf{O}=\mathrm{Norm}(\mathbf{Q}(\mathbf{K}^{\top}\mathbf{V})).$ (3)
@@ 3.2 LASP
LASP tiles sequences over the cluster. Follow the thought-of-tiling, LASP partitions the input sequences into multiple sub-sequence chunks, distributing these chunks individually across different GPUs. For linear attention in a casual setting, in order to fully exploit the advantage of right-product in linear attention, we categorize the attention computation for chunks into two distinct types: intra-chunks and inter-chunks. Intra-chunks involve conventional attention computation, while inter-chunks leverage the kernel tricks associated with linear attention’s right-product. Further details regarding the intricate mechanisms of LASP in data distribution, forward pass, and backward pass are expounded upon below. A visualization of LASP is presented in Fig. 1.
Figure 1: Visualization of LASP. Left: The chunk-level linear attention computation with a causal mask can be segmented into two categories: intra-chunk and inter-chunk computations. Intra-chunk computations, corresponding to the diagonal elements (in diagonal orange boxes) of the mask matrix, utilize traditional left-product methods. While inter-chunk computations, corresponding to the lower triangular boxes, employ efficient right-product methods for computation. Right: This panel illustrates the P2P communication mechanism employed by LASP. The input sequence $\mathbf{X}$ is divided into multiple sub-sequence chunks $\{\cdots,\mathbf{X}_{i},\mathbf{X}_{i+1},\cdots\}$, each processed by different model instances across distinct devices. For each device $i$, $Q_{i}$, $K_{i}$, and $V_{i}$ are computed from its respective input chunk $\mathbf{X}_{i}$. Notably, the communication operations between devices are designed to be complementary in the forward and backward passes. Specifically, in the forward pass, $\mathbf{KV}$ matrices are sent from device $i$ to device $(i+1)$, and in the backward pass, $\mathbf{dKV}$ matrices are sent back from device $(i+1)$ to device $i$.
@@
Figure 2: LASP Data Distribution. Left: An example of data distribution with two input sequences and eight GPUs. Right: Complete data distribution algorithm.
Data Distribution.
For an input sequence of length $N$, we establish its embedding space representation denoted as $\mathbf{X}\in\mathbb{R}^{N\times d}$ with a feature dimension of $d$. In the LASP framework, $\mathbf{X}$ is evenly partitioned into $T$ chunks, where $T$ is called the sequence parallel size, which must be divisible by the distributed world size $W$. These segmented data chunks are subsequently assigned to the respective GPUs. It is essential to note that different sequence parallel groups receive dissimilar data batches. However, within the same group, all data chunks originate from an identical batch of data. A comprehensive depiction of the data distribution process in LASP is provided in Algorithm 1.
Additionally, an illustrative example of data distribution in LASP is presented in Fig. 2, where the distributed world size is characterized by $W=8$, the sequence parallel size by $T=4$, the number of sequence parallel groups by $G=2$, and the sequence parallel source rank list by $R_{{src}}=[0,4]$. For the first batch Seq0, the input sequence ${\mathbf{X}}$ undergoes partitioning into $T$ chunks $\{\mathbf{X}_{1},\mathbf{X}_{2},...,\mathbf{X}_{T}\}$ along the sequence dimension, subsequently transmitted to the first rank in SP-Group0, which corresponds to global rank 00. The data chunks on global rank 0 are then scattered to global ranks $\{0,1,2,3\}$ within SP-Group0, where each rank only retains a single chunk. The subsequent batch Seq1 follows a similar manner, being assigned to global ranks $\{4,5,6,7\}$ within SP-Group1.
Forward Pass.
@@
13:     Save $\mathbf{KV}_{{t-1}}$ as $\mathbf{KV}_{i}$ for backward computation.
@@
15:     Compute $\mathbf{O}_{t}=\mathbf{O}_{\mathrm{t,intra}}+\mathbf{O}_{\mathrm{t,inter}}$.
@@
To streamline derivations, the $\mathrm{Norm}(\cdot)$ operator in Eq. (2) is temporarily omitted. Additionally, we consider a normal case where $W=T$, indicating $G=W/T=1$. In this scenario, GPU with rank 00 consolidates all split sub-sequences in a batch, subsequently distributing them to all GPUs across the entire distributed world. It is noteworthy that the scenario where the sequence parallel size is not equal to world size is discussed in Sec.3.5.
We first define $\mathbf{kv}$ and $\mathbf{KV}$ as the intermediate memory state vector and matrix, respectively. Without loss of generality, we add $\lambda$ as the decay rate in linear attention with casual masking, choosing $\lambda=1$ yields the ordinary linear attention (Qin et al., 2024a; Sun et al., 2023). In the forward pass of linear attention computation with casual masking, the $s$-th output can be calculated as
$\mathbf{o}_{s}^{\top}=\mathbf{q}_{s}^{\top}\sum_{i\leq s}\lambda^{s-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}.$ (4)
Rewrite in a recurrence form, we have
$\mathbf{kv}_{0}=$ $0\in\mathbb{R}^{d\times d},\quad\mathbf{kv}_{s}=$ $\lambda\mathbf{kv}_{s-1}+\mathbf{k}_{s}\mathbf{v}_{s}^{\top},\quad\mathbf{o}_{s}^{\top}=\mathbf{q}_{s}^{\top}(\mathbf{kv}_{s}),$ (5)
@@
$\mathbf{kv}_{s}=\sum_{i\leq s}\lambda^{s-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}$ (6)
is the activation memory state in the forward pass with $s$-th input.
In SP, given data chunk $\mathbf{X}_{t}$ on rank $i$, the query, key and value corresponding to $\mathbf{X}_{t}$ is $\mathbf{Q}_{t}=\mathbf{X}_{t}\mathbf{W}_{Q}$, $\mathbf{K}_{t}=\mathbf{X}_{t}\mathbf{W}_{K}$, $\mathbf{V}_{t}=\mathbf{X}_{t}\mathbf{W}_{V}$. Note that we assume $T=W$ here, their indices are thus equivalent, i.e., $t=i$. The output within the $t$-th chunk can be calculated as
$\mathbf{O}_{\mathrm{t,intra}}=[(\mathbf{Q}_{t}\mathbf{K}_{t}^{\top})\odot\mathbf{M}]\mathbf{V}_{t}.$ (7)
@@
$\mathbf{o}_{s+C}^{\top}=$ $\mathbf{q}_{s+C}^{\top}\sum_{i\leq s+C}\lambda^{s+C-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}=$ $\mathbf{q}_{s+C}^{\top}\sum_{i=C+1}^{C+s}\lambda^{s+C-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}$ $+\lambda^{s}\mathbf{q}_{s+C}^{\top}\sum_{i\leq C}\lambda^{C-i}\mathbf{k}_{i}\mathbf{v}_{i}^{\top}.$ (8)
The resulted first part in Eq. (8) corresponds to the computation on previous chunks, and the second part corresponds to the computation on the current chunk. In SP, Eq. (8) can be rewritten in the chunk form as
$\mathbf{O}_{\mathrm{t,inter}}=\bm{\Lambda}\mathbf{Q}_{t}\mathbf{K}\mathbf{V}_{t-1},$ (9)
where $\mathbf{KV}_{t}=\mathbf{kv}_{tC}$. Note that the calculation of the inter-chunk for the $t$-th chunk depends on the activation state of previous $(t-1)$ chunk, i.e., $\mathbf{KV}_{t-1}$, which is calculated on rank $(i-1)$. Thus a P2P communication operation Recv should be performed to pull $\mathbf{KV}_{t-1}$ from rank $(i-1)$ to rank $i$. Then the activation state $\mathbf{KV}_{t}$ should be updated for subsequent inter-chunk attention computation at $(t+1)$-th chunk. The update rule of $\mathbf{KV}_{t}$ at $t$-th chunk is
$\mathbf{KV}_{t}$ $=\sum_{s\leq tC}\lambda^{tC-s}\mathbf{k}_{s}\mathbf{v}_{s}^{\top}=\lambda^{C}\!\!\!\!\sum_{s\leq(t-1)C}\!\!\!\!\lambda^{(t-1)C-s}\mathbf{k}_{s}\mathbf{v}_{s}^{\top}+\!\!\!\!\!\!\sum_{s=(t-1)C+1}^{tC}\!\!\!\!\!\!\lambda^{tC-s}\mathbf{k}_{s}\mathbf{v}_{s}^{\top}$ (10)
$=\lambda^{C}\mathbf{KV}_{t-1}+\left(\mathrm{diag}\{\lambda^{C-1},\ldots,1\}\mathbf{K}_{t}\right)^{\top}\mathbf{V}_{t}=\lambda^{C}\mathbf{KV}_{t-1}+\left(\lambda^{C}\mathbf{\Lambda}^{-1}\mathbf{K}_{t}\right)^{\top}\mathbf{V}_{t}.$